Operations Report – January 19th

Hi Mission Support!

Crew 203 Operations Report 19-JAN-2019

SOL: 7

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 17hr 45min; Turned on last night (18Jan2019) at 16:29; Turned off this morning (19Jan2019) at 9:15; Turned on tonight (19Jan2019) at 19:33

Solar SOC – Turned on (18Jan2019) 91%; Turned off (19Jan2019) 100%; Turned on (19Jan2019) 60%

Diesel Reading – Just going under 1/2 tank.

Propane Reading – 23%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – Over 60%; About 325 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 65%; 194 gallons

Water (loft) – At level marker 9

Static to Loft Pump used – Yes; At 9:45 to refill tank

Water Meter: 01402506

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for EVA to Tank Wash

Hours: 68.8

Beginning charge: 100%

Ending charge: 52%

Currently charging: Yes

Opportunity rover used: Not used

Hours: 45.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, used for EVA to Tank Wash

Hours: 70.6

Beginning charge: 100%

Ending charge: 65%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to go to Grand Junction to take Shannon to the doctor.

CrewCar used and why, where? Off campus

General notes and comments: We receive a new toilet brush and three new spare batteries too.

Summary of internet: Nothing to report

Summary of suits and radios: Suit #3 still pending for review.

Summary of Hab operations: A broken toilet brush was replaced. Battery of smoke detector was replaced.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Battery of smoke detector was replaced.

Summary of RAM operations: Battery of smoke detector was replaced.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report

A little tired due to today’s EVA but we keep fighting 😉